OpenStudy (anonymous):

What is restriced domain? ..I have to give the restricted domain of (f*g) and (g*f)

OpenStudy (anonymous):

Well, normally this comes up for inverse functions. If you have a function like x^2 and you invert it as is, it will not be one to one. You "restrict the domain" to the part that will be 1 to 1 when inverted.

OpenStudy (anonymous):

So what's the restriced domain of (f*g) and (g*f) ... the star being that little circle

OpenStudy (anonymous):

This question is incomplete.

OpenStudy (anonymous):

ok what if it said Let f(x)=x^2 and g(x)=1/x-5

OpenStudy (anonymous):

That is vital information. Can you identify what f(g(x)) and g(f(x)) are??